# Production of dileptonic bound states in the Higgs boson decay

### v2

@@ I Introduction
After the discovery of the Higgs boson, a period of more detailed study of its properties and decay characteristics through various channels began. Such theoretical research now relies on existing experimental data, the collection of which is steadily expanding [1]. Plans have begun to emerge for various future accelerators in which there will be a significant increase in the production of Higgs bosons, and, therefore, the possibility of more accurately determining various parameters of the Higgs sector will open up [2, 3]. As often happens when formulating programs for future colliders, the task is set to search for events that do not fit into the scheme of the Standard Model. Among the rare decays of the Higgs boson, one can distinguish a class of processes in which bound states of fundamental particles arise, such as mesons and baryons [4, 5, 6, 7]. The theoretical interest in these reactions is due to the fact that, among other things, they can additionally test models for describing bound states of quarks. The number of such works has grown significantly over the past 10 years, and the methods for calculating the observed characteristics have received significant development [8]. Another group of reactions for the formation of bound states of particles consists of those reactions in which bound states of leptons (positronium, dimuonium, ditauonium) can be produced. In reactions of this type there are no uncertainties caused by the nonperturbative interaction of quarks. Various processes of production and decay of positronium already have a rich history, which included critical periods of divergence between theory and experiment. Other bound states of leptons have been studied much less. Thus, it can be noted that calculations of the characteristics of dimuonium have been carried out quite a long time ago with great accuracy, but the bound state itself has not yet been observed [9]. Various concepts of dimuonium and ditauonium production are discussed in many papers [10, 11, 12]. One possibility for the production of lepton bound states is connected with the decay of the Higgs boson. In this work, we study the processes of both single and pair production of dimuonium and ditauonium within the framework of relativistic method of describing the production of bound states of particles. This method was used previously to calculate the production of charmonium and $B_{c}$ mesons [13, 14, 15]. It allows one to obtain values of the decay width taking into account relativistic effects. It is useful to note that the observation of orthodimuonium or orthoditauonium is connected with subsequent decay into three photons.
@@ II Production of single leptonic bound states
Let us first consider the single production of lepton bound states in the decay of $H\to\gamma+\bar{l}l$, $H\to Z+\bar{l}l$. The Higgs boson-lepton vertex is determined by the factor [16]:
@@
The lepton-photon (lepton-Z-boson) mechanism of single production of dimuonium and ditauonium is determined by three amplitudes presented in Fig. 1(a,b,c). The initial expression for the amplitude of such a process can be represented as a convolution of the amplitude of the production of a photon and a lepton-antilepton pair with the Bethe-Salpeter wave function, which describes, within the framework of the four-dimensional formalism, a lepton-antilepton bound state moving with 4-momentum $P$. Transforming the amplitude during the transition to three-dimensional formalism and using the law of transformation of the quasipotential wave function during the transition from the moving frame of reference to the rest frame of the lepton bound state (see, [17, 18]), we obtain two decay amplitudes (Fig. 1(a,b)) in the form:
@@
where $r$ is the Higgs boson four-momentum, $\varepsilon_{\gamma}$ is the photon polarization vector, $\varepsilon_{l\bar{l}}$ is the polarization vector of dimuonium (ditauonium), $k$ is the photon four-momentum, $\varepsilon(p)=\sqrt{p^{2}+m^{2}}$, $v=P/M$, $M$ is the mass of $l\bar{l}$ state. $p$ is the lepton relative four-momentum. The expressions for the amplitudes (2), (3) are obtained as a result of the Lorentz transformations in the same way as in our works [13, 14, 15]. They make it possible to take into account relativistic corrections connected with the momentum of relative motion of leptons $p$. The 4-momenta of the photon and leptonium (bound state lepton-anti-lepton) are given as arguments in the amplitudes (2), (3), although it should be remembered that they lie on the mass surface, and the following conservation law is satisfied: $r=k+P$.
@@
The studies we carried out in [13, 14] show that it is necessary to take into account the loop ($W$-boson and quark) production mechanism $Z+(l\bar{l})$ shown in Fig. 4. Although formally an additional degree of $\alpha$ appears in it compared to the $ZZ$-boson mechanism, the fine structure constant is not a decisive factor in determining the order of magnitude of the contribution. It is useful to mention here that a loop mechanism with W-boson and quark loops is needed to describe the decay process $H\to Z+\gamma$, the first evidence for which was presented recently in [19]. Numerous particle mass ratio parameters begin to play a more significant role in loop production mechanism. The loops of heavy quarks and the $W$-boson make their contribution to the structure functions $g_{1}$ and $g_{2}$, which can be represented as follows:
@@
The lepton-photon decay mechanism in Fig. 1 gives a decay width of fifth order in $\alpha$. But our analysis shows that the lepton-photon mechanism in Fig. 2 has a significantly larger numerical contribution, although it has a higher order $\alpha^{7}$. This occurs due to simultaneously changing mass coefficients in the amplitude of such a decay. We previously studied a similar decay mechanism in connection with the production of heavy quarkonia in [13, 14, 15].
@@
where $t=\frac{M_{H}^{2}}{4m_{Q}^{2}}$ or $t=\frac{M_{H}^{2}}{4m_{W}^{2}}$, $m_{W}$ is the mass of $W$-boson, $m_{Q}$ is the mass of heavy quark in the quark loop. The structure functions $A_{Q,W}(t)$, $B_{Q,W}(t)$ can be obtained using an explicit expression for a loop integrals [20, 21, 22]. In the case where one photon is real, the only contribution comes from the structure function $A(t)$ which satisfies the dispersion relation:
@@
Imaginary part $ImA(t)$ can be calculated using the Mandelstam-Cutkosky rule [23]:
@@ III.1 Lepton - photon production mechanism A
Taking into account the law of transformation of the wave function of a bound state upon transition to a rest frame, we obtain wave functions in (26) in the form [13, 14]:
@@
The calculation of trace in (26) of the product of $\gamma$-factors and various simplifications are performed in the Form package [24]. Considering small factors $\frac{{\bf p}}{m}$, $\frac{{\bf q}}{m}$, we perform an expansion on them in the numerator of the amplitude, while preserving the second-order terms. Then the numerator of this amplitude takes the form:
@@ IV.1 Relativistic parameters in production amplitude
To calculate relativistic corrections to the $H$-boson decay amplitude, we use the same approach as in our previous works [13, 14]. After calculating the trace in the numerator of the amplitudes, corrections of a certain order of smallness in the relative momentum of leptons in the bound state are identified. We take into account second-order corrections in ${\bf p}^{2}/m^{2}$, ${\bf q}^{2}/m^{2}$ and introduce an expansion of the form:
@@
The parameters $\omega_{i}$ for $i>1$ are proportional to higher powers of $\alpha$ than (68) and are omitted from now on. The results (68) are obtained using the following expression [25, 26]:
@@ IV.2 Relativistic corrections to the wave function
Wave function (65) was obtained from Schrödinger equation solution with Coulomb potential. In our previous works we had estimated relativistic correction to wave function by accounting corresponding terms in quark-antiquark potential [13, 14, 15]. In present research we also can estimate leading order relativistic correction to wave function as follows [9, 27, 28, 29, 30]:
[…]
@@
It is useful to note that the same expression for the correction can be obtained using the relativistic correction to the wave function from [27]:
@@
[…]
@@
[…]
@@ V Numerical results and Conclusions
In Table 1 we present the numerical results of calculating the decay widths of the Higgs boson with the production of single lepton bound states and a photon or Z-boson. The obtained values for $\Gamma$ of these rare decays are small, and their observation is possible only at high-luminosity colliders being designed in the future, when the production of a significant number of Higgs bosons ($\sim 10^{10})$ will be possible [2, 3]. Taking into account the main channel of the decay of lepton ortho-states into three photons, one can expect their reliable detection.
@@
Final state Result of work [8] Our result
$(e^{+}e^{-})+\gamma$ $3.5\times 10^{-12}$ $0.88\times 10^{-10}$
@@
In calculating decay widths, we examine the contribution of various decay mechanisms, the relative contribution of which changes in the transition from positronium to dimuonium and ditauonium. Numerous mass parameters $r_{i}$ in the presented analytical formulas for decay widths, which determine the final result, change greatly during the transition between lepton bound states. But we can definitely say that the contribution from W-boson and quark loops (see Fig. 7,8) has the greatest values, along with the contribution of the ZZ-mechanism in the case of leptonium production. The widths of rare Higgs boson decays were calculated earlier in [8]. In general, we can say that our results and [8] are consistent, although there are some differences in individual decay widths. These differences may be related to the different production mechanisms of lepton bound states investigated in our work.
@@ Acknowledgements.
The work is supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS” (Grant No. 22-1-1-23-1).

### v3

@@ I Introduction
After the discovery of the Higgs boson, a period of more detailed study of its properties and decay characteristics through various channels began. Such theoretical research now relies on existing experimental data, the collection of which is steadily expanding [higgs]. Plans have begun to emerge for various future accelerators in which there will be a significant increase in the production of Higgs bosons, and, therefore, the possibility of more accurately determining various parameters of the Higgs sector will open up [clic, fcc]. As often happens when formulating programs for future colliders, the task is set to search for events that do not fit into the scheme of the Standard Model. Among the rare decays of the Higgs boson, one can distinguish a class of processes in which bound states of fundamental particles arise, such as mesons and baryons [p1, p2, p3, p4]. The theoretical interest in these reactions is due to the fact that, among other things, they can additionally test models for describing bound states of quarks. The number of such works has grown significantly over the past 10 years, and the methods for calculating the observed characteristics have received significant development [higgs1]. Another group of reactions for the formation of bound states of particles consists of those reactions in which bound states of leptons (positronium, dimuonium, ditauonium) can be produced. In reactions of this type there are no uncertainties caused by the nonperturbative interaction of quarks. Various processes of production and decay of positronium already have a rich history, which included critical periods of divergence between theory and experiment. Other bound states of leptons have been studied much less. Thus, it can be noted that calculations of the characteristics of dimuonium have been carried out quite a long time ago with great accuracy, but the bound state itself has not yet been observed [sgk1]. Various concepts of dimuonium and ditauonium production are discussed in many papers [higgs2, higgs3, enterria]. One possibility for the production of lepton bound states is connected with the decay of the Higgs boson. In this work, we study the processes of both single and pair production of dimuonium and ditauonium within the framework of relativistic method of describing the production of bound states of particles. This method was used previously to calculate the production of charmonium and $B_{c}$ mesons [apm2021, apm2023, apm2023a]. It allows one to obtain values of the decay width taking into account relativistic effects. It is useful to note that the observation of orthodimuonium or orthoditauonium is connected with subsequent decay into three photons.
@@ II Production of single leptonic bound states
Let us first consider the single production of lepton bound states in the decay of $H\to\gamma+\bar{l}l$, $H\to Z+\bar{l}l$. The Higgs boson-lepton vertex is determined by the factor [roma]:
@@
The lepton-photon (lepton-Z-boson) mechanism of single production of dimuonium and ditauonium is determined by three amplitudes presented in Fig. 1(a,b,c). The initial expression for the amplitude of such a process can be represented as a convolution of the amplitude of the production of a photon and a lepton-antilepton pair with the Bethe-Salpeter wave function, which describes, within the framework of the four-dimensional formalism, a lepton-antilepton bound state moving with 4-momentum $P$. Transforming the amplitude during the transition to three-dimensional formalism and using the law of transformation of the quasipotential wave function during the transition from the moving frame of reference to the rest frame of the lepton bound state (see, [faustov, brodsky]), we obtain two decay amplitudes (Fig. 1(a,b)) in the form:
@@
where $r$ is the Higgs boson four-momentum, $\varepsilon_{\gamma}$ is the photon polarization vector, $\varepsilon_{l\bar{l}}$ is the polarization vector of dimuonium (ditauonium), $k$ is the photon four-momentum, $\varepsilon(p)=\sqrt{p^{2}+m^{2}}$, $v=P/M$, $M$ is the mass of $l\bar{l}$ state. $p$ is the lepton relative four-momentum. The expressions for the amplitudes (2), (3) are obtained as a result of the Lorentz transformations in the same way as in our works [apm2021, apm2023, apm2023a]. They make it possible to take into account relativistic corrections connected with the momentum of relative motion of leptons $p$. The 4-momenta of the photon and leptonium (bound state lepton-anti-lepton) are given as arguments in the amplitudes (2), (3), although it should be remembered that they lie on the mass surface, and the following conservation law is satisfied: $r=k+P$.
@@
The studies we carried out in [apm2021, apm2023] show that it is necessary to take into account the loop ($W$-boson and quark) production mechanism $Z+(l\bar{l})$ shown in Fig. 4. Although formally an additional degree of $\alpha$ appears in it compared to the $ZZ$-boson mechanism, the fine structure constant is not a decisive factor in determining the order of magnitude of the contribution. It is useful to mention here that a loop mechanism with W-boson and quark loops is needed to describe the decay process $H\to Z+\gamma$, the first evidence for which was presented recently in [Aad]. Numerous particle mass ratio parameters begin to play a more significant role in loop production mechanism. The loops of heavy quarks and the $W$-boson make their contribution to the structure functions $g_{1}$ and $g_{2}$, which can be represented as follows:
@@
The lepton-photon decay mechanism in Fig. 1 gives a decay width of fifth order in $\alpha$. But our analysis shows that the lepton-photon mechanism in Fig. 2 has a significantly larger numerical contribution, although it has a higher order $\alpha^{7}$. This occurs due to simultaneously changing mass coefficients in the amplitude of such a decay. We previously studied a similar decay mechanism in connection with the production of heavy quarkonia in [apm2021, apm2023, apm2023a].
@@
where $t=\frac{M_{H}^{2}}{4m_{Q}^{2}}$ or $t=\frac{M_{H}^{2}}{4m_{W}^{2}}$, $m_{W}$ is the mass of $W$-boson, $m_{Q}$ is the mass of heavy quark in the quark loop. The structure functions $A_{Q,W}(t)$, $B_{Q,W}(t)$ can be obtained using an explicit expression for a loop integrals [pp1, pp2, pp3]. In the case where one photon is real, the only contribution comes from the structure function $A(t)$ which satisfies the dispersion relation:
@@
Imaginary part $ImA(t)$ can be calculated using the Mandelstam-Cutkosky rule [t4]:
@@ III.1 Lepton - photon production mechanism A
Taking into account the law of transformation of the wave function of a bound state upon transition to a rest frame, we obtain wave functions in (26) in the form [apm2021, apm2023]:
@@
The calculation of trace in (26) of the product of $\gamma$-factors and various simplifications are performed in the Form package [form]. Considering small factors $\frac{{\bf p}}{m}$, $\frac{{\bf q}}{m}$, we perform an expansion on them in the numerator of the amplitude, while preserving the second-order terms. Then the numerator of this amplitude takes the form:
@@ IV.1 Relativistic parameters in production amplitude
To calculate relativistic corrections to the $H$-boson decay amplitude, we use the same approach as in our previous works [apm2021, apm2023]. After calculating the trace in the numerator of the amplitudes, corrections of a certain order of smallness in the relative momentum of leptons in the bound state are identified. We take into account second-order corrections in ${\bf p}^{2}/m^{2}$, ${\bf q}^{2}/m^{2}$ and introduce an expansion of the form:
@@
The parameters $\omega_{i}$ for $i>1$ are proportional to higher powers of $\alpha$ than (68) and are omitted from now on. The results (68) are obtained using the following expression [ibk1994, apm1998]:
@@ IV.2 Relativistic corrections to the wave function
Wave function (65) was obtained from Schrödinger equation solution with Coulomb potential. In our previous works we had estimated relativistic correction to wave function by accounting corresponding terms in quark-antiquark potential [apm2021, apm2023, apm2023a]. In present research we also can estimate leading order relativistic correction to wave function as follows [sgk1, lepage1, lepage2, khriplovich, penin]:
$\psi^{(1)}_{1S}(0)=\int{\tilde{G}}_{1S}(0,{\bf r})\Delta V({\bf r})\psi^{(0)}_{1S}({\bf r})d{\bf r},$ (70)
@@
It is useful to note that the same expression for the correction can be obtained using the relativistic correction to the wave function from [lepage1]:
@@
$\Delta\psi_{rel}^{(3)}=\psi^{C}_{1S}(0)\frac{2W^{2}}{m^{2}\pi^{4}}\int R({\bf p},{\bf q})d{\bf p}d{\bf q}\left[\frac{2}{p}\arctan(p)-\frac{1}{p^{2}+1}\right]=$
@@
$\Delta V_{vp}(r)=-\frac{\alpha^{2}}{3\pi}\int_{1}^{\infty}\rho(\xi)d\xi\frac{1}{r}e^{-2m_{e}\xi r},~~~\rho(\xi)=\frac{\sqrt{\xi^{2}-1}(2\xi^{2}+1)}{\xi^{4}}.$ (91)
@@ V Numerical results and Conclusions
In Table 1 we present the numerical results of calculating the decay widths of the Higgs boson with the production of single lepton bound states and a photon or Z-boson. The obtained values for $\Gamma$ of these rare decays are small, and their observation is possible only at high-luminosity colliders being designed in the future, when the production of a significant number of Higgs bosons ($\sim 10^{10})$ will be possible [clic, fcc]. Taking into account the main channel of the decay of lepton ortho-states into three photons, one can expect their reliable detection.
@@
Final state Result of work [higgs1] Our result
$(e^{+}e^{-})+\gamma$ $3.5\times 10^{-12}$ $1.10\times 10^{-11}$
@@
In calculating decay widths, we examine the contribution of various decay mechanisms, the relative contribution of which changes in the transition from positronium to dimuonium and ditauonium. Numerous mass parameters $r_{i}$ in the presented analytical formulas for decay widths, which determine the final result, change greatly during the transition between lepton bound states. But we can definitely say that the contribution from W-boson and quark loops (see Fig. 7,8) has the greatest values, along with the contribution of the ZZ-mechanism in the case of leptonium production. The widths of rare Higgs boson decays were calculated earlier in [higgs1]. In general, we can say that our results and [higgs1] are consistent, although there are some differences in individual decay widths. These differences may be related to the different production mechanisms of lepton bound states investigated in our work.
@@ Acknowledgements.
The work is supported by the Foundation for the Advancement of Theoretical Physics and Mathematics “BASIS” (Grant No. 22-1-1-23-1). The authors are grateful to D. d’Enterria and V.D. Le for useful discussions.